how many numbers have an absolute value of 5/4th's??? What are they? How did you get this???

Answer by jim_thompson5910(35256)

5/4 and -5/4 both have the the absolute value of 5/4

ie

|5/4| = 5/4
|-5/4| = 5/4

Both numbers -5/4 and 5/4 have this same absolute value because they are both 5/4 units away from the number 0.

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim